OpenStudy (anonymous):

2/5 is greater than or less than 2/5^2

OpenStudy (dangerousjesse):

Less than, of course..

OpenStudy (anonymous):

just testing

OpenStudy (anonymous):

wait what? its greater no?

OpenStudy (jdoe0001):

eheh

OpenStudy (dangerousjesse):

Nope..

OpenStudy (jdoe0001):

how many 25th in the 2/5?

OpenStudy (jdoe0001):

if you want to convert the 2/5 to 25ths just multiply top and bottom by 5

OpenStudy (dangerousjesse):

\(\frac{2}{5}<\frac{2}{5}^2\) is true..

OpenStudy (anonymous):

i got it wrong it was greater than

OpenStudy (jdoe0001):

\(\bf \cfrac{2}{5}\qquad \cfrac{2}{5^2} \\ \quad \\ \cfrac{2}{5}\qquad \cfrac{2}{25} \\ \quad \\ \cfrac{2\times 5}{5\times 5}\qquad \cfrac{2}{25}\implies ?\)

OpenStudy (jdoe0001):

so 2/5 is really when converting it to 25ths is \(\bf \cfrac{2}{5}\qquad \cfrac{2}{5^2} \\ \quad \\ \cfrac{2}{5}\qquad \cfrac{2}{25} \\ \quad \\ \cfrac{2\times 5}{5\times 5}\qquad \cfrac{2}{25}\implies \cfrac{10}{25} > \cfrac{2}{25}\)

OpenStudy (dangerousjesse):

Oh! so it's 5 to the power of two, not 2/5 to the power of two. I read that wrong.